# MoReMouse: Monocular Reconstruction of Laboratory Mouse

### v1

@@ Abstract
Laboratory mice play a crucial role in biomedical research, yet accurate 3D mouse surface motion reconstruction remains challenging due to their complex non-rigid geometric deformations and textureless appearance. Moreover, the absence of structured 3D datasets severely hinders the progress beyond sparse keypoint tracking. To narrow the gap, we present MoReMouse, the first monocular dense 3D reconstruction network tailored for laboratory mice. To achieve this goal, we highlight three key designs. First, we construct the first high-fidelity dense-view synthetic dataset for mice, by rendering our self-designed realistic Gaussian mouse avatar. Second, MoReMouse adopts a transformer-based feedforward architecture with triplane representation, achieving high-quality 3D surface generation from a single image. Third, we create geodesic-based continuous correspondence embeddings on mouse surface, which serve as strong semantic priors to improve reconstruction stability and surface consistency. Extensive quantitative and qualitative experiments demonstrate that MoReMouse significantly outperforms existing open-source methods in accuracy and robustness. Video results are available at https://zyyw-eric.github.io/MoreMouse-webpage/.
Figure 1: We present MoReMouse, the first monocular dense 3D reconstruction framework for laboratory mice. Given a single-view input (top-left, captured using an iPhone 15 Pro), MoReMouse predicts high-fidelity surface geometry and appearance using a transformer-based triplane architecture. The model is trained on a dataset constructed from real multi-view videos fitted to a Gaussian mouse avatar. Our method supports fast inference and outputs RGB renderings, semantic embeddings, and normal maps from a single image.
@@ 1 Introduction
Laboratory mice, as widely used model animals, are fundamental to disease modeling and neuroscience studies. Understanding their three-dimensional (3D) motion dynamics provides valuable insights for medical and life sciences[13].
@@
Over the last decade, deep learning has revolutionized both 2D and 3D animal pose estimation, with methods such as DeepLabCut [24], SLEAP [29] and DANNCE [11] becoming standard tools across diverse species including mice. However, these methods are limited to sparse keypoint tracking and ignore important 3D surface motion. Although several papers have tried to build articulated mouse mesh models [6, 4, 1] and use them for mouse surface tracking [2, 17], they still face two key drawbacks. First, manually designed skinning weights are infeasible to represent the complex dynamic deformations of a mouse, resulting in severe self-penetration during its free movement. Second, existing mouse mesh fitting methods require heavy and sensitive multi-view optimization process, hindering its application for more convenient single view scenarios.
The rapid evolution of 3D generative AI has blurred the boundary between 3D reconstruction and 3D generation [39], where large scale 3D datasets such as Objaverse [10, 9] play a crucial role in this trend. Thanks to the datasets, Large Reconstruction Models (LRMs) [14, 39, 38] demonstrate single-view feedforward 3D reconstruction for general objects, significantly outperform traditional feedback approaches like DreamFusion [30] in both reconstruction quality and efficiency. Although LRMs have the potential to revolutionize monocular mouse reconstruction, they require ground truth 3D geometries or dense-view images for supervision, which is unavailable for laboratory mice. Additionally, mice exhibit highly dynamic non-rigid deformations, and their furry skins lack prominent visual features, making it more challenging to reconstruct surface motion of mice.
@@
In this paper, we present MoReMouse, the first monocular dense 3D reconstruction model for laboratory mouse. MoReMouse is built upon transformer architectures and triplane representation, enabling efficient, feedforward 3D reconstruction. To compensate for the scarcity of dense-view 3D mouse dataset, we construct the first animatable Gaussian avatar of mouse using an articulated mouse model [6] and Gaussian Splatting [16]. Inspired by recent human avatar modeling [20], we propose a 2D UV-based intermediate mouse pose representation to drive Gaussian point clouds. Thanks for the free-view rendering ability and diverse pose space of the avatar, we build a synthetic 64-view dataset, enabling effective training of MoReMouse network.
@@
Despite of the synthetic dataset, the framework of MoReMouse integrates multiple architectural and training enhancements inspired by LRMs. Specifically, we employ a two-stage training strategy, initially pretraining the model with triplane-NeRF [7, 25] representations, followed by fine-tuning with triplane-DMTet[35] for higher-fidelity reconstructions. To further improve the stability and robustness of dense surface reconstruction, we introduce geodesic-based feature embeddings, which enhance semantic consistency of surfaces across frames. In experiments, we compare MoReMouse with existing open-sourced feedforward large reconstruction models such as TripoSR [39], Triplane-GS [50], InstantMesh [45] and LGM [38] on our synthetic dataset, proving the superiority of MoReMouse for monocular mouse reconstruction and novel view synthesis. Qualitative results on real-world videos captured by hand-held IPhone further confirm the generalization ability of MoReMouse (Fig. 1). As a whole, MoReMouse represents a crucial step toward high-fidelity, efficient 3D reconstruction of mouse, and will facilitate broader applications in biomedical research. To summarize, our key contributions include:
@@
  • We introduce MoReMouse, the first monocular dense 3D reconstruction model for laboratory mouse, surpassing existing open-source alternatives.
@@
  • We develop the first Gaussian avatar model for mice and generate a large scale dense-view synthetic mouse dataset for training and evaluation.
@@
  • We integrate geodesic-based continuous correspondence embeddings with feedforward 3D reconstruction, improving single-view reconstruction stability and quality.
@@ 2 Related work
Previously, most 3D mouse motion capture toolboxes reconstruct mouse 3D keypoints by algebraic triangulation [27, 15, 12, 26], voxel-based triangulation [11] or even by markers [22, 23]. However, 3D keypoints provide only sparse geometric structures, neglecting dense surface information which is essential to represent highly dynamic mouse body shapes or interpret social touch information [2, 17]. To represent the mouse surface,  [5] proposes an articulated mesh model based on computed tomography scans. Although this model is useful for the motion capture of mouse surface from multiview videos [2], its manually designed skinning weight could not capture the realistic dynamic surface motion of a moving mouse. ArMo [4] presents a similar mouse mesh model, but lacks realistic geometric details. For the rat, which shares a shape similar to that of the mouse, researchers pay similar efforts to build articulated mesh model [1] yet face the same challenge. In this paper, we try to overcome the limitation of explicit articulated representation of mouse by reconstructing the implicit surface of it.
2.2 Monocular Reconstruction of Animals
@@
Since the proposal of SMAL (Skinned Multi-Animal Linear) model which represents several quadrupedal species [52], predicting the pose and shape parameters of SMAL from images has dominated the monocular reconstruction of animals [53, 54, 3, 44, 32, 33, 21]. Although they have demonstrated satisfactory performance for species such as dogs, horses, cattle, and tigers, they cannot reconstruct rodents due to the limited shape space of SMAL. The AWOL model [51] further extends SMAL to represent rodents using CLIP [31] and supports monocular shape estimation. However, it still relies on a predefined articulated model, without the ability to express geometric deformation beyond linear blend skinning. Some recent papers try to build animal avatars from videos by augmenting the skinning weights or adding additional deformations. For example, BANMo [47] and RAC [48] build fully implicit cat and dog avatars using neural deformation field and differentiable rendering [25], Gart [18] and Animal Avatars [34] build dog avatars with SMAL geometric embedding and neural appearance. However, these methods all require a heavy optimization process during inference, and rely on dense pose estimation such as CSE [28] or SMAL parameters for initialization.
2.3 Large Reconstruction Model
Large reconstruction models have recently achieved substantial progress in generating high-quality 3D content [14, 19, 41, 46, 50, 49, 36, 37]. These models are typically trained in an end-to-end fashion to predict implicit 3D representations, often through triplane-based neural radiance fields. Some approaches operate from a single input image [14, 41, 46, 50, 49], while others use multi-view images generated by novel view synthesis techniques [19, 36]. Meanwhile, UNet-based architectures have also proven effective in generating multiplane or Gaussian representations [42, 38], providing alternative and efficient formats for 3D reconstruction.
However, these methods generally depend on access to 3D supervision, such as ground-truth shapes or high-quality multi-view data, during training. This reliance poses a significant challenge for in-the-wild applications, where diverse and scalable 3D datasets remain scarce compared to their 2D counterparts [8, 10, 43]. In addition, their reconstructions often exhibit limitations in geometric detail and texture quality, making them less suitable for scenarios that demand high visual fidelity.
@@ 3 Method
Our goal is to achieve dense geometry reconstruction of a mouse from a single viewpoint via a feed-forward 3D generative approach, where representation and supervision play crucial roles in ensuring high-quality generation.
To construct dataset for effective supervision (Sec. 3.1), we design the first Gaussian avatar for mouse (Sec. 3.1.2), which introduces an additional bias on top of the mesh model to better capture complex mouse surface deformations from videos (Sec. 3.1.3).
Our proposed MoReMouse model builds upon an LRM-like structure (Sec. 3.2) and exploits the strengths of both Triplane-NeRF for an initial coarse training phase and Triplane-DMTet for fine-grained training (Sec. 3.2.1). Furthermore, to enhance tracking stability and preserve surface consistency, we introduce a geodesic-distance-based feature embedding (Sec. 3.2.2). These embeddings serve as additional semantic priors, improving the robustness of dense reconstruction. The details of our training and optimization procedure are provided in (Sec. 3.2.3).
3.1 Synthetic Dataset using Gaussian Avatar
Figure 2: Visualization of the articulated mouse mesh and the learned geodesic embedding. Left: our mesh model consists of 140 articulated joints and 14,522 vertices, forming a kinematically consistent 3D structure for pose fitting. Right: color-coded visualization of the learned geodesic feature embedding, where principal components are mapped to HSV color space to enhance surface separability and promote reconstruction consistency.
3.1.1 Mouse Parametric Model
As shown in Fig 2, we employ the established mouse mesh model used in  [2] as avatar control signals. The model consists of $K=140$ articulated joints and 14,522 vertices, and conforms to Linear Blend Skinning (LBS) for pose control. The whole parameter set is denoted as $\Psi$, including joint rotations $\theta\in\mathbb{R}^{3K}$, global translation $t\in\mathbb{R}^{3}$, global rotation $r\in\mathbb{R}^{3}$, global scale $s\in\mathbb{R}$ and bone length deformation parameters $B\in\mathbb{R}^{21}$ (Please refer to  [2] for more details). We decompose $\Psi$ into local parameters $\Psi_{l}=\{\theta,B\}$ and global parameters $\Psi_{g}=\{r,t,s\}$ for the sake of simplicity.
3.1.2 Gaussian Mouse Avatar
Figure 3: Pipeline of the Gaussian Mouse Avatar Generation. The process begins with Mesh Fitting, followed by Gaussian model training.
Gaussian point clouds [16] use an explicit parameterization for 3D representation. Each Gaussian point is characterized by a mean position $\mu\in\mathbb{R}^{3}$ and an anisotropic covariance matrix $\Sigma$:
@@ 3.1.2 Gaussian Mouse Avatar
Inspired by Animatable Gaussians [20], we employ texture-based deformation transfer to train a neural representation of the Gaussian point cloud avatar, incorporating bias correction beyond the skeletal model, as shown in Fig. 3. Our key idea is to parameterize the mouse UV texture, mapping each valid texel in the UV space to a corresponding Gaussian point, where the texture color represents the Gaussian’s position.
To obtain the canonical space representation, we set $\Psi_{g}=0$ and drive the mesh model using only $\Psi_{l}$, resulting in canonical vertex coordinates $V_{\Psi_{l}}$. The UV-mapped positions of these vertices are used to interpolate Gaussian locations in the UV texture via barycentric coordinates. This defines the baseline Gaussian position map $\mu_{\Psi_{l}}$:
@@
During Gaussian optimization, we incorporate a CNN-based offset prediction module. A convolutional neural network StyleUNet [40] predicts the position offset map $\Delta\mu_{\Psi_{l}}$, conditioned on the UV coordinate map $\mu_{\Psi_{l}}$, to refine the Gaussian positions:
$\mu^{\prime}_{\Psi_{l}}=\mathcal{D}_{\Psi_{l}}(\mu_{0}+\Delta\mu_{\Psi_{l}}).$ (3)
The CNN also estimates additional Gaussian attributes, including color $c_{\Psi_{l}}$, opacity $o_{\Psi_{l}}$, covariance rotation parameters $q_{\Psi_{l}}$, and scaling parameters $s_{\Psi_{l}}$. Then the Gaussian Splatting rendering function generates the view-dependent rendered image $I_{c}^{\text{rend}}$:
@@
3.1.3 Avatar Training
To train the Gaussian avatar, we define loss functions between the real image $I_{c}$ and the rendered image $I_{c}^{\text{rend}}$:
@@ 3.1.3 Avatar Training
where $\mathcal{L}_{\text{SSIM}}$ quantifies structural similarity, while $\mathcal{L}_{\text{LPIPS}}$ assesses perceptual similarity. The loss weights are set to $\lambda_{\text{SSIM}}=0.2$ and $\lambda_{\text{LPIPS}}=0.1$. Additionally, we enforce Total Variation (TV) loss to regularize Gaussian point cloud smoothness.
We use a 6-view mouse video sequence ”markerless_mouse_1” proposed in [11] for avatar training. Each sequence captures 18,000 frames of a single freely moving mouse in an open field at 1152x1024@100FPS, where the mouse mesh models are provided by  [2]. Note that, we only use 800 frames for training, which are uniformly sampled from the first 8000 frames of ”markerless_mouse_1”. Avatar training takes 400k steps.
3.2 MoReMouse Architecture
3.2.1 Transformer and Triplane Representation
Similar to LRM [14], MoReMouse adopts a Transformer and triplane-based architecture, designed for monocular 3D reconstruction. The model consists of three key components: an image encoder, an image-to-triplane decoder, and a triplane-based geometric representation that integrates both NeRF [25] and DMTet [35] for implicit and explicit geometry modeling.
First, the image encoder employs a pre-trained DINOv1 backbone to extract a latent feature vector from the input image. Then, a large Transformer decoder employs cross-attention to project these image features into a 3D triplane representation. This triplane-based geometric representation encodes color, density, and additional spatial attributes at any queried 3D location.
During the initial training phase, to regularize the convergence of the spatial density field, we do not introduce explicit geometry. Instead, we employ NeRF for differentiable volumetric rendering, allowing gradients to propagate throughout the entire space.
Once the density field has sufficiently converged, we introduce DMTet as an explicit geometric representation for further refinement. DMTet provides well-defined surfaces and enables direct geometric supervision.

### v2

@@ Abstract
Laboratory mice, particularly the C57BL/6 strain, are essential animal models in biomedical research. However, accurate 3D surface motion reconstruction of mice remains a significant challenge due to their complex non-rigid deformations, textureless fur-covered surfaces, and the lack of realistic 3D mesh models. Moreover, existing visual datasets for mice reconstruction only contain sparse viewpoints without 3D geometries. To fill the gap, we introduce MoReMouse, the first monocular dense 3D reconstruction network specifically designed for C57BL/6 mice. To achieve high-fidelity 3D reconstructions, we present three key innovations. First, we create the first high-fidelity, dense-view synthetic dataset for C57BL/6 mice by rendering a realistic, anatomically accurate Gaussian mouse avatar. Second, MoReMouse leverages a transformer-based feedforward architecture combined with triplane representation, enabling high-quality 3D surface generation from a single image, optimized for the intricacies of small animal morphology. Third, we propose geodesic-based continuous correspondence embeddings on the mouse surface, which serve as strong semantic priors, improving surface consistency and reconstruction stability, especially in highly dynamic regions like limbs and tail. Through extensive quantitative and qualitative evaluations, we demonstrate that MoReMouse significantly outperforms existing open-source methods in both accuracy and robustness.
Project pagehttps://zyyw-eric.github.io/MoreMouse-webpage/
Figure 1: We present MoReMouse, the first monocular dense 3D reconstruction framework for laboratory mice. Given a single-view input (top-left, captured with an iPhone 15 Pro), MoReMouse predicts high-fidelity surface geometry and appearance within 0.9 seconds using a transformer-based triplane architecture (middle). To assist model training, we render a dense-view synthetic dataset by building the first Gaussian mouse avatar from sparse-view real videos (bottom-left). Our method outputs RGB renderings, semantic embeddings, and normal maps from a single image (right). Corresponding video results are provided in the supplementary material.
@@ 1 Introduction
Laboratory mice are widely used model organisms in biomedical research, especially for disease modeling, genetic studies, and neuroscience. Among them, the C57BL/6 strain is the most commonly used inbred line due to its well-characterized genome and relevance in neurodegenerative, oncological, and immunological research. Accurate 3D motion reconstruction of C57BL/6 mice is essential for behavioral phenotyping.
@@
Over the last decade, deep learning has revolutionized both 2D and 3D animal pose estimation, with methods such as DeepLabCut (mathis2018deeplabcut), SLEAP (pereira2022sleap) and DANNCE (dunn2021geometric) becoming standard tools across diverse species including mice. However, these methods are limited to sparse keypoint tracking and ignore important 3D surface motion. To this end, template-based papers have tried to build articulated mouse mesh models (49_bolanos2021three) and used them for mouse 3D tracking (an2023three). However, the manually designed skinning weights are infeasible to represent the complex dynamic deformations of a mouse, resulting in severe self-penetration during its free movement. Moreover, mesh fitting methods require heavy and sensitive multi-view optimization, hindering its application for more convenient single view scenarios. Template-free animal reconstruction methods such as 3D-Fauna (li2024learning) lacks the knowledge of mouse-specific appearance and geometry, therefore has difficulty in recovering 3D C57BL/6 mice.
The rapid evolution of 3D generative AI has blurred the boundary between 3D reconstruction and 3D generation, with large-scale datasets such as Objaverse (deitke2023objaverse) playing a crucial role in this trend. Fueled by such data, Large Reconstruction Models (LRMs) (hong2023lrm; tochilkin2024triposr; tang2024lgm) have demonstrated high efficiency and accuracy in single-view, feedforward 3D reconstruction of general objects. More recent generative frameworks, such as Trellis (xiang2025structured) based on rectified flow, and CLAY (zhang2024clay) based on diffusion, further push the scalability and controllability of 3D generation by learning powerful structured latent representations. Although these models show great promise, they do not perform well on laboratory mice due to the lack of large-scale 3D mice supervision or dense-view mice images. Moreover, as these methods only considers single-image cases, no temporally semantic correspondences are ensured for long term dynamic object reconstruction.
@@
In this paper, we present MoReMouse, the first monocular dense 3D reconstruction model specifically designed for C57BL/6 laboratory mouse. MoReMouse is built upon transformer architectures and triplane representation, enabling efficient, feedforward 3D reconstruction, as shown in Fig. 1. To compensate for the scarcity of dense-view 3D mouse dataset, we construct the first animatable Gaussian avatar of mouse (AGAM for short) using an articulated mouse model (49_bolanos2021three) and Gaussian Splatting (kerbl2023gaussian). Inspired by recent human avatar modeling (li2024animatable), we use a 2D UV-based intermediate mouse pose representation to drive Gaussian point clouds. Thanks for the free-view rendering ability and diverse pose space of the avatar, we build a synthetic 64-view dataset, enabling effective training of MoReMouse network.
@@
Based on the new synthetic dataset, the framework of MoReMouse integrates multiple architectural and training enhancements inspired by LRMs. Specifically, we employ a two-stage training strategy, initially pretraining the model with triplane-NeRF (chan2022efficient; mildenhall2021nerf) representations, followed by fine-tuning with triplane-DMTet(shen2021deep) for higher-fidelity reconstructions. To further improve the stability and robustness, we introduce geodesic-based feature embeddings, which enhance semantic consistency of surfaces across frames.
We evaluate MoReMouse on both synthetic and real-world datasets. The synthetic benchmark is rendered using our AGAM, while the real-world evaluation uses a 4-view video sequence of single freely moving C57BL/6 mouse captured by PointGray industrial cameras. Results demonstrate the superior performance of MoReMouse in monocular mouse reconstruction and novel view synthesis. As a whole, MoReMouse represents a significant step toward practical and high-fidelity monocular 3D reconstruction of C57BL/6 mice, with potential to support broader research in computational modeling and analysis of small animals. To summarize, our key contributions include:
@@
  • We introduce MoReMouse, the first monocular dense 3D reconstruction model tailored for laboratory mice, specifically the C57BL/6 strain, and show its superiority over existing open-source alternatives.
@@
  • We develop the first animatable Gaussian avatar model for mice and use it to generate a large-scale, dense-view synthetic dataset for training and evaluation.
@@
  • We integrate geodesic-based continuous correspondence embeddings into feedforward reconstruction, improving single-view surface stability and detail preservation.
@@ 2 Related work
Previously, most methods reconstruct mouse 3D keypoints by algebraic triangulation (35_nath2019using; 36_karashchuk2021anipose; han2024multi; monsees2022estimation), voxel-based triangulation (dunn2021geometric) or even by optical markers (marshall2021pair). However, 3D keypoints provide only sparse geometric structures, neglecting essential dense surface information. To represent the mouse surface,  (49_bolanos2021three) proposes an articulated mesh model based on computed tomography scans, and is used for multi-view mesh fitting (an2023three). ArMo (bohnslav2023armo) presents a similar mouse mesh model, but lacks true anatomical bones. For rats, which are similar to mice in shape, similar mesh models are built (aldarondo2024virtual; klibaite2025mapping). All these mesh models are limited to the manually designed skinning weights, which could not capture the realistic dynamic surface motion of mice.
2 Monocular Reconstruction of Animals
@@
Since the proposal of SMAL model which represents several quadrupedal species (63_zuffi20173d), predicting the pose and shape parameters of SMAL from images has dominated the monocular reconstruction of animals (zuffi2019three; biggs2020wldo; xu2023animal3d; rueegg2022barc; lyu2025animer). Although they perform well for quadrupeds, they cannot reconstruct rodents due to the limited shape space of SMAL. The AWOL model (zuffi2024awol) extends SMAL to represent rodents using CLIP (radford2021learning), yet its pose and shape realism is still limited. Some recent papers try to build animal avatars from videos by augmenting the skinning weights or adding additional deformations, such as BANMo (yang2022banmo), Gart (lei2024gart) or AnimalAvatars (Sabathier2024AnimalAR). However, these methods all require a heavy optimization process during inference, and rely on dense pose estimation or accurate SMAL parameters for initialization. Template-free methods such as MagicPony (wu2023magicpony) and 3D Fauna (li2024learning) leverage collections of single-view images of deformable species captured in the wild to jointly learn canonical models and pose regressors. However, these approaches do not generalize well to mice due to domain and geometry gaps.
3 Large Reconstruction Model
Large reconstruction models have recently achieved substantial progress in generating high-quality 3D content (hong2023lrm; Li2023Instant3DFT; zou2024triplane; gslrm2024; sun2024dreamcraft3d++). These models are typically trained in an end-to-end fashion to predict implicit 3D representations, often through triplane-based neural radiance fields. Recent methods improve quality and efficiency by learning generative models over compact latent codes derived via VAE-like encoders (hu2025simulating). Some focus purely on shape (zhang20233dshape2vecset; li2024craftsman3d), while others incorporate appearance (gupta20233dgen; xiong2024octfusion). CLAY (zhang2024clay) presents a unified system for geometry and PBR texture generation. Trellis (xiang2025structured) introduces structured latent codes via rectified flow transformers, jointly modeling geometry and appearance for versatile decoding into NeRFs, Gaussians, and meshes. However, these methods generally depend on access to 3D supervision (deitke2023objaverse; wu2023omniobject3d), such as ground-truth shapes or high-quality multi-view data, during training. The data scarcity is one of the main challenges for training an LRM for laboratory mice.
@@ 3 Method
Our goal is to reconstruct dense surface geometry of a laboratory mouse from a single image in a feedforward manner. We focus specifically on the C57BL/6 strain, a widely used model organism in biomedical research. We first introduce our Gaussian mouse avatar (AGAM) and synthetic dataset in Sec. 1, then the structure of MoReMouse in Sec. 2.
1 Synthetic Dataset using Gaussian Avatar
Figure 2: Left: the articulated mouse mesh we used. Right: color-coded geodesic feature embedding, promoting surface separability and reconstruction consistency.
Mouse Parametric Model
As shown in Fig 2, we modify the mouse model from  (49_bolanos2021three) following the approach of  (an2023three), where redundant mesh vertices (teeth and tone) are removed. The model consists of $N_{J}=140$ articulated joints and $N_{V}=13059$ vertices, and conforms to Linear Blend Skinning (LBS) for posing. Poses of this refined model serve as the avatar control signals. The whole parameter set is denoted as $\Psi$, including joint rotations $\theta\in\mathbb{R}^{3N_{J}}$, global translation $t\in\mathbb{R}^{3}$, global rotation $r\in\mathbb{R}^{3}$, global scale $s\in\mathbb{R}$ and bone length deformation parameters $B\in\mathbb{R}^{21}$ (Please refer to  (an2023three) for more details). We decompose $\Psi$ into local parameters $\Psi_{l}=\{\theta,B\}$ and global parameters $\Psi_{g}=\{r,t,s\}$ for the sake of simplicity.
Gaussian Mouse Avatar
Figure 3: Pipeline of the Gaussian Mouse Avatar Generation. The process begins with Mesh Fitting, followed by Gaussian Avatar training. “CNN” here is StyleUNet.
Gaussian point clouds (kerbl2023gaussian) use an explicit parameterization for 3D representation. Each Gaussian point is characterized by a mean position $\mu\in\mathbb{R}^{3}$ and an anisotropic covariance matrix $\Sigma$:
@@ 1 Synthetic Dataset using Gaussian Avatar
The pipeline of our AGAM is shown in Fig. 3. Our key idea is to parameterize the mouse UV texture, mapping each valid texel in the UV space to a corresponding Gaussian point, where the texture color represents the Gaussian’s position. To achieve global-pose invariant controlling of AGAM, we set $\Psi_{g}=0$ and drive the mesh model using only $\Psi_{l}$, resulting in canonical vertex coordinates $V_{\Psi_{l}}$. The UV-mapped positions of these vertices are used to interpolate Gaussian locations via barycentric coordinates. This defines the baseline Gaussian position map $\mu_{\Psi_{l}}$:
@@
The core network of AGAM is a set of StyleUNets (wang2023styleavatar) that predict the position offset map $\Delta\mu_{\Psi_{l}}$ and other Gaussian attributes $c_{\Psi_{l}}$, $o_{\Psi_{l}}$, $q_{\Psi_{l}}$, and $s_{\Psi_{l}}$ conditioned on the UV coordinate map $\mu_{\Psi_{l}}$. Note that, the Gaussian positions $\mu^{\prime}_{\Psi_{l}}$ are derived by $\mu^{\prime}_{\Psi_{l}}=\mathcal{D}_{\Psi_{l}}(\mu_{0}+\Delta\mu_{\Psi_{l}}).$ Then the Gaussian Splatting rendering function generates the view-dependent rendered image $I_{c}^{\text{rend}}$:
@@
Avatar Training
We define loss functions between the real image $I_{c}$ and the rendered image $I_{c}^{\text{rend}}$ as:
@@
where $\mathcal{L}_{\text{SSIM}}$ is structural similarity loss, $\mathcal{L}_{\text{LPIPS}}$ is perceptual similarity loss, $\lambda_{\text{SSIM}}=0.2$ and $\lambda_{\text{LPIPS}}=0.1$. We use Total Variation (TV) loss to enforce point smoothness. See the supplementary material for further details.
Since C57BL/6 mice usually show similar texture, we only use a 6-view mouse video sequence “markerless_mouse_1” proposed by (dunn2021geometric) for avatar training. This sequence captures 18,000 frames of a single freely moving mouse in an open field at 1152x1024@100FPS, where the mouse mesh models are borrowed from  (an2023three). Note that, we only use 800 frames for training, which are uniformly sampled from the first 8000 frames of “markerless_mouse_1”. Avatar training takes 400k steps.
Dense-view Dataset
We drive AGAM using the poses from all frames of the “markerless_mouse_1” video sequence, and render evenly distributed 64-view synthetic images. Among the full 18,000 frames, we use the first 6,000 frames for training and the last 6,000 frames for testing. Note that, the 800 frames used for avatar creation all belong to the first 6000 frames, therefore the poses in testset are unseen. To construct the training set, we sample two sets of 64 viewpoints per frame, rendering 12,000 multi-view scenes in total. The optical axes of these viewpoints point at the scene origin. Please refer to the supplementary file for more information about view selection and data samples.
2 MoReMouse Architecture
Transformer and Triplane Representation
Similar to LRM (hong2023lrm), MoReMouse adopts a Transformer and triplane-based architecture, designed for monocular 3D reconstruction. As illustrated in Fig. 4, the model consists of three key components: an image encoder, an image-to-triplane decoder, and a triplane-based geometric representation that integrates both NeRF (mildenhall2021nerf) and DMTet (shen2021deep) for implicit and explicit geometry modeling.
We use a pretrained DINOv2 encoder to extract latent features from the input image. These features are decoded via a transformer with cross-attention to generate a triplane representation, which encodes color, density, and semantic attributes for any 3D query point.
During the initial training stage, we employ NeRF for volumetric rendering without introducing explicit geometry, enabling smooth gradient propagation through the entire density field. After convergence, we switch to DMTet to obtain an explicit surface representation for fine-level geometric supervision.